OpenStudy (anonymous):

A student opens a mathematics book to two facing pages. The product of the page numbers is 272. Find the page numbers

OpenStudy (anonymous):

The pages are 16 and 17. 16 * 17 = 272.

OpenStudy (anonymous):

let number on one page be x then the number on the succeeding/facing page will be x+1 ATQ, x (x+1) = 272 x^2 - x - 272 = 0 If you solve this quadratic equation u will get your answer...... if u need further help, pls let me know....

OpenStudy (anonymous):

16 and 17

OpenStudy (anonymous):

correct.... I made a small mistake in the equation... it shud read x^2 + x - 272 = 0

OpenStudy (anonymous):

Also harkirat your solution is incorrect. Your quadratic equation would wield a value of 17 for x and therefore the other page would be 18, which multiplies to 306. page a = x page b = x+1 A better solution is x(x+1) = 272 x^2 + x = 272 x^2+x-272=0 and so x = 16 x+1 = 17

OpenStudy (anonymous):

@genius12 I never gave any solutions, I had just formed the equation......... I hv also realised the mistake and told Cali about it .............☺